synthetic devision?
Use synthetic division and the Remainder Theorem to evaluate P(C). P(x) = x3 + 2x2 – 7, c = -2 P(-2) =
Answer #1

Solution Given - P(x)= x + 2x²- 7 C= -2 -2/12 -7 h ő - Remainder P(-2) = (-2)² + 2 (-2) ²7 = -8+8-7 = 0-7 [P(-2) = -7 Am